OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2) 11 −11 1 −1

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

Using the distributive property of: a(b+c) = ab+ac You get: 3x + 12 = 5x - 10 You want to isolate x on one side, so subtract 3x from both sides to get: 12 = 2x - 10 Now solve for x by adding 10 to both sides and dividing by 2: x = 22/2 = 11 So x = 11 Derp, sorry, subtracted wrong xD